OpenStudy (anonymous):

What is the common ratio of this geometric sequence 1 50 2 30 3 18 4 10.8 5 6.48

OpenStudy (anonymous):

Answers ->> A. -20 B. 0.6 C. 1.67 D. 20

OpenStudy (ybarrap):

30/50 = ? 18/30 = ? 10.8/18 = ? etc Each of these should be equal and is the answer to this question

OpenStudy (anonymous):

What

OpenStudy (anonymous):

Oh it's 0.6?